OpenStudy (anonymous):

5 square root of 6x end root plus 8 square root of 14 x end root minus 9 square root of 6x end root of 18 square root of 14 x

OpenStudy (goformit100):

@marcoduuuh

OpenStudy (jackellyn):

\[5\sqrt{6x}+8\sqrt{14x}-9\sqrt{6x}+18\sqrt{14x}\] is that the question?